# Decaying turbulence beneath surface waves

### v1

@@ Abstract
This paper explores decaying turbulence beneath surface waves that is initially isotropic and shear-free. We start by presenting phenomenology revealed by wave-averaged numerical simulations: an accumulation of angular momentum in coherent vortices, suppression of kinetic energy dissipation, and the development of depth-alternating jets. We interpret these features through an analogy with rotating turbulence (Holm, 1996), wherein the curl of the Stokes drift, ×$\nabla\times\bm{u}^{\mathrm{S}}$, takes on the role of the background vorticity (for example, $(f_{0}+\beta y)\hskip 1.00006pt\bm{\hat{z}}$ on the beta plane). We pursue this thread further by showing that a two-equation model proposed by Bardina et al. (1985) for rotating turbulence reproduces the simulated evolution of volume-integrated kinetic energy. This success of the two-equation model — which explicitly parameterizes wave-driven suppression of kinetic energy dissipation — carries implications for modeling turbulent mixing in the ocean surface boundary layer. We conclude with a discussion about a wave-averaged analogue of the Rossby number appearing in the two-equation model, which we term the “pseudovorticity number” after the pseudovorticity ×$\nabla\times\bm{u}^{S}$. The pseudovorticity number is related to the Langmuir number in an integral sense.
@@ 1 Introduction
Surface waves enhance near-surface ocean turbulent mixing, especially in summertime and in the tropics where boundary layers are often shallow, sunny, windy, and wavy Sullivan & McWilliams (2010). Turbulence driven by surface wind stress and affected by surface waves is usually called “Langmuir turbulence” (McWilliams et al., 1997), implying a connection between wave-catalyzed turbulent coherent structures and the structure of a laminar, wave-catalyzed shear instability that Craik & Leibovich (1976) called “Langmuir circulation”.
@@
We interpret the results of these simulations in section 2, leveraging an analogy between wave-averaged and rotating dynamics first proposed by Holm (1996). Like rotation, surface waves catalyze an inverse cascade — without exchanging energy with the flow — and cause angular momentum to accumulate in coherent vortices, while suppressing kinetic energy dissipation. We observe the development of alternating jets, as in beta-plane turbulence. In section 3 we further the analogy by adapting a phenomenological model proposed by Bardina et al. (1985) for decaying rotating turbulence to the wave-averaged case, showing that the model also describes the decay of volume-averaged kinetic energy beneath surface waves. This model suggests a new way to incorporate wave effects into two-equation models (for example, Harcourt 2015) by explicitly representing the surface-wave-driven suppression of kinetic energy dissipation.
@@ 2 Simulations of decaying turbulence beneath surface waves
The incompressible Euler momentum equation in the presence of a background vorticity $\bm{\Omega}$,
@@
describes both rotating flows and wave-averaged flows beneath steady, horizontally-uniform surface wave fields (Craik & Leibovich, 1976; Suzuki & Fox-Kemper, 2016; Wagner et al., 2021). In the wave-averaged case, $\bm{u}$ in (1) is is the Lagrangian-mean momentum and $p$ is Eulerian-mean kinematic pressure. To use (1) for rotating flows we take
@@
where $f$ is the Coriolis parameter and $\hskip 1.00006pt\bm{\hat{z}}$ is the axis of rotation. For wave-averaged flows we take
@@
where $\bm{u}^{\mathrm{S}}=u^{\mathrm{S}}(z)\hskip 1.00006pt\bm{\hat{x}}$ is the Stokes drift of a surface wave field propagating in the $\hskip 1.00006pt\bm{\hat{x}}$-direction. In the context of equation (1), the main difference between rotating turbulence and turbulence beneath steady surface waves is the spatial structure of $\partial_{z}u^{\mathrm{S}}$: in shallow water $\partial_{z}u^{\mathrm{S}}\sim z$ to leading-order, while $\partial_{z}u^{\mathrm{S}}$ decays exponentially for deep water waves.
@@ 2.1 Coherent structures in rotating and wave-averaged turbulence
To illustrate the similarity of rotating turbulence and turbulence beneath surface waves we conduct large eddy simulations of equation (1) in a unit cube with three choices for $\bm{\Omega}$,
@@
We produce an initial condition for the three simulations by conducting a preliminary simulation initialized with kinetic energy spectrum
@@
The initial simulation is run for a duration $t_{0}$ until $t=0$, defined as the time when the mean-square vorticity has decayed to $\omega_{\text{rms}}(t=0)=10$. The velocity field is then saved to disk to be used as an initial condition in subsequent runs starting from $t=0$. The simulations are conducted with Oceananigans (Ramadhan et al., 2020; Wagner et al., 2025), which discretizes (1) with a finite volume method. We use a nominally 9th-order Weighted, Essentially Non-Oscillatory advection scheme (Shu, 2020), which dissipates kinetic energy at the grid scale. Scripts that reproduce simulations in this paper are stored on GitHub; see the Data availability statement.
We focus first on the evolution of the relative vorticity $\bm{\omega}$ defined in equation (6). Figure 1 shows vorticity components for the three cases after $t=1000$ time units: figure 1(a) shows vertical vorticity $\zeta$, while while figures 1(b) and (c) show the horizontal vorticity $\eta$. Figure 1(a) and (b) for rotating and wave-affected turbulence, respectively, both exhibit the formation of coherent structures and relatively greater vorticity levels than the unorganized, small amplitude isotropic vorticity in figure 1(c). Figure 2 is similar, except that figure 2(a) shows $\zeta$ in rotating turbulence in the $xy$-plane, while figures 2(b) and (c) show $\eta$ in the $xz$ plane for wave-averaged and isotropic turbulence, respectively.
@@ 2.2 Zonation and the analogy of wave-averaged turbulence with beta-plane turbulence
Note that turbulence beneath shallow water waves is homeomorphic to $\beta$-plane turbulence Rhines (1975) — with $\bm{\Omega}_{\text{$\beta$-plane}}=\beta y\hskip 1.00006pt\bm{\hat{z}}$ — modulo on the $xz$ plane instead of on the $xy$ plane. One of the most striking results of this similarity is the propensity for turbulence beneath surface waves to develop “zonal jets” — coherent, alternating jets in the direction of surface wave propagation (and perpendicular to the pseudovorticity direction). This similarity allows us, therefore, to borrow intuition on structure formation in $\beta$-plane turbulence (for example, Huang & Robinson (1998); Farrell & Ioannou (2007); Srinivasan & Young (2012); Constantinou et al. (2014); for an overview see Constantinou (2015); Farrell & Ioannou (2019); Marston & Tobias (2023).)
@@
To illustrate this, we consider two additional cases with pseudovorticity
@@
$\bm{\Omega}_{\mathrm{deep}}=\tfrac{1}{4}\mathrm{e}^{8(z-1)}\hskip 1.00006pt\bm{\hat{y}}\,.\qquad\text{and}\qquad\bm{\Omega}_{\mathrm{weak}}=\tfrac{1}{8}z\hskip 1.00006pt\bm{\hat{y}}\,.$ (7)
@@
These and subsequent simulations use $384^{3}$ finite volume cells.
@@
Figure 3 shows time-series of the $y$-momentum $v$, which plays the role that vertical velocity plays in rotating turbulence, as well as vertical profiles of the horizontally-averaged $x$-momentum,
$U(z,t)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\int u\,\mathrm{d}x\,\mathrm{d}y\,,$ (8)
@@
The profiles of $U$ on the right side of figure 3 exhibit the development of depth-alternating jets, which to our knowledge has not yet been observed in shear-free wave-modified turbulence. Comparing the slices of $v$ for medium and weak waves reveals how the medium-strong waves induce a strong inverse cascade, more coherent vortices, and fewer small-scale motions than the weak waves.
@@
Figure 3: The evolution of cross-wave momentum $v$ in the $xz$-plane (6 panels on the left) and horizontally-averaged along-wave-momentum $u$ (2 panels on the right) at $t=40,400$ and for three wave fields: “deep” (left panel, red lines), “medium” (middle panel, orange lines), and “weak” (right panel, blue lines). Dashed lines in the top right plot show the Stokes drift profile (normalized) for each case. The $u$ profiles and light gray “zero lines” are spaced apart by the scale $\delta u=10^{-2}$. The $u$-profiles at right show the development of depth-alternating jets, including a counter-wave surface jet for all cases. The $v$-slices are similar between the three cases at early times, but at later times exhibit strong, localized wave-impacts in their respective regions of significant Stokes shear. Note that $v$ plays the role that vertical velocity plays in rotating turbulence.
@@ 3 Phenomenological model for the evolution of kinetic energy
Figure 4 plots time-series of the normalized kinetic energy $k(t)/k(t=0)$ for the three cases presented in figure 1, as well as three additional cases that use $\bm{\Omega}=Sz\hskip 1.00006pt\bm{\hat{y}}$ with $S=(1/2,1/8,1/16)$. Figure 4 illustrates another common feature to rotating turbulence and turbulence beneath surface waves: a suppression of the kinetic energy dissipation rate, such that at long times the kinetic energy levels off.
Inspired by Bardina et al. (1985), we model $k(t)$ with the phenomenological two-equation system
@@
The free parameter $a$ may be constrained by considering isotropic turbulence with $\Omega=0$. In this case we expect the turbulent kinetic energy to decay according to $k\sim t^{6/5}$ (Saffman, 1967), which implies $\frac{\mathrm{d}}{\mathrm{d}t}k=-\frac{6}{5}\frac{k}{t}$, and thus in turn, via (10), leads to
$\epsilon=-\frac{6}{5}\frac{k}{t}\,,\quad\text{so that}\quad\frac{\mathrm{d}}{\mathrm{d}t}\epsilon=-\frac{11}{5}\frac{k}{t^{2}}\,.$ (12)
@@
[…]
@@
$\log\frac{\epsilon}{\epsilon_{0}}=a\log\left(\frac{k}{k_{0}}\mathrm{e}^{-b\Omega t}\right)\,,$ (14)
where $\epsilon_{0}$ and $k_{0}$ are the dissipation and kinetic energy at $t=0$. Inserting (14) into (10), integrating in time and rearranging then produces the solution
$\frac{k}{k_{0}}=k_{0}\left\{1+\frac{1}{nb\Omega}\frac{\epsilon_{0}}{k_{0}}\left(1-\mathrm{e}^{-b\Omega t}\right)\right\}^{-n}\,,$ (15)
where $n=1/(a-1)=6/5$. Taking the limit $\Omega\to 0$ (or $t\to 0$ with finite $\Omega$), we obtain the corresponding solution for isotropic turbulence,
@@
Figure 4: The decay of kinetic energy $k(t)$ in isotropic, rotating, and surface-wave-modulated turbulence. Solid lines show kinetic energy normalized by it’s initial value, $k/k_{0}$, computed from large eddy simulations. Dashed lines show $k/k_{0}$ given by (15), which solves the phenomenological two-equation system in (10)–(11). The initial dissipation rate $\epsilon_{0}$ in (15) is computed numerically from $k(t)$. We use $a=11/6$ and estimate $b$ from (18). For the rotating case, $\Omega=1/4$ yields $b=0.033$. For the four surface-wave-modulated cases we use $b=0.036$, where $\Omega\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\int\partial_{z}u^{\mathrm{S}}{\,\mathrm{d}}z=\left(1/4,1/8,1/16,1/32\right)$. Note that using $a=1.75$ along with commensurate adjustments to $b$ matches the simulation data even more closely.

### v2

# Phenomenology of decaying turbulence beneath surface waves

@@ Abstract
This paper explores decaying turbulence beneath surface waves that is initially isotropic and shear-free. We start by presenting phenomenology revealed by wave-averaged numerical simulations: an accumulation of angular momentum in coherent vortices perpendicular to the direction of wave propagation, suppression of kinetic energy dissipation, and the development of depth-alternating jets. We interpret these features through an analogy with rotating turbulence (Holm, 1996), wherein the curl of the Stokes drift, ×$\bm{\nabla}\times\bm{u}^{\mathrm{S}}$, takes on the role of the background vorticity (for example, $(f_{0}+\beta y)\hskip 1.00006pt\bm{\hat{z}}$ on the beta plane). We pursue this thread further by showing that a two-equation model proposed by Bardina et al. (1985) for rotating turbulence reproduces the simulated evolution of volume-integrated kinetic energy. This success of the two-equation model — which explicitly parametrizes wave-driven suppression of kinetic energy dissipation — carries implications for modeling turbulent mixing in the ocean surface boundary layer. We conclude with a discussion about a wave-averaged analogue of the Rossby number appearing in the two-equation model, which we term the “pseudovorticity number” after the pseudovorticity ×$\bm{\nabla}\times\bm{u}^{S}$. The pseudovorticity number is related to the Langmuir number in an integral sense.
@@ 1 Introduction
Surface waves enhance near-surface ocean turbulent mixing, especially in summertime and in the tropics where boundary layers are often shallow, sunny, windy, and wavy (Sullivan & McWilliams, 2010). Turbulence driven by surface wind stress and affected by surface waves is usually called “Langmuir turbulence” (McWilliams et al., 1997), implying a connection between wave-catalyzed turbulent coherent structures and the structure of a laminar, wave-catalyzed shear instability that Craik & Leibovich (1976) called “Langmuir circulation”.
@@
We interpret the results of these simulations in section 2, leveraging an analogy between wave-averaged and rotating dynamics first proposed by Holm (1996). Like rotation, surface waves catalyze a transfer of energy from smaller to larger scales (hereby referred to as ‘inverse cascade’) within the turbulent flow beneath the waves — without exchanging energy with the flow — and cause angular momentum to accumulate in coherent vortices, while suppressing kinetic energy dissipation. This contrasts our case with the interaction between balanced flows and internal waves, wherein non-catalytic interactions can occur even for steady wave fields (Thomas, 2023). Moreover, unlike in shear-driven Langmuir turbulence where coherent structures tend to align with the direction of the Lagrangian-mean shear (Sullivan & McWilliams, 2010; Van Roekel et al., 2012), the coherent vortices that we observe forming in the absence of shear are instead oriented perpendicular to the direction of surface wave propagation. We also observe the development of alternating jets, as in beta-plane turbulence.
In section 3 we further the analogy by adapting a phenomenological model proposed by Bardina et al. (1985) for decaying rotating turbulence to the wave-averaged case, showing that the model also describes the decay of volume-averaged kinetic energy beneath surface waves. This model suggests a new way to incorporate wave effects into two-equation models (for example, Harcourt (2015)) by explicitly representing the surface-wave-driven suppression of kinetic energy dissipation.
@@ 2 Simulations of decaying turbulence beneath surface waves
The incompressible, inviscid momentum equation in the presence of a background vorticity $\bm{\Omega}$,
@@
describes both rotating flows without surface waves and also wave-averaged flows beneath steady, horizontally-uniform surface-wave fields (see for example Craik & Leibovich, 1976; Suzuki & Fox-Kemper, 2016; Holm, 1996; Wagner et al., 2021). In the wave-averaged case, $p$ and $\bm{u}$ in (1) represent the Eulerian-mean kinematic pressure and Lagrangian-mean velocity, respectively. The Lagrangian-mean velocity is the “total” velocity responsible for the advection of mass, momentum, tracers, and particles, and may be decomposed into an Eulerian-mean component plus “Stokes drift” correction (van den Bremer & Breivik, 2018; Vanneste & Young, 2022, see for example). The wave-averaged momentum equation may be derived by asymptotic expansion leveraging a time-scale separation between rapidly oscillating surface waves and the slower evolution of $\bm{u}$ (Craik & Leibovich, 1976; Leibovich, 1980; Vanneste & Young, 2022). The appearance of surface waves in (1) as an effective Coriolis force is discussed by Holm (1996).
For rotating flows, the background vorticity $\bm{\Omega}$ in (1) takes the form
@@
where $f$ is the Coriolis parameter and $\hskip 1.00006pt\bm{\hat{z}}$ is the axis of rotation. For wave-averaged flows beneath a uniform surface wave field propagating in the $x$ direction, the background vorticity $\bm{\Omega}$ is alternatively
@@
where $\bm{u}^{\mathrm{S}}=u^{\mathrm{S}}(z)\hskip 1.00006pt\bm{\hat{x}}$ is the surface wave Stokes drift (Craik & Leibovich, 1976; Suzuki & Fox-Kemper, 2016; van den Bremer & Breivik, 2018; Vanneste & Young, 2022). We refer to ×$\bm{\nabla}\times\bm{u}^{\mathrm{S}}$ — the curl of the surface-wave pseudomomentum $\bm{u}^{\mathrm{S}}$ (Andrews & McIntyre, 1978) — as the “pseudovorticity”. In the context of equation (1), the main difference between rotating turbulence and turbulence beneath steady surface waves is the spatial structure of $\partial_{z}u^{\mathrm{S}}$: in shallow water $\partial_{z}u^{\mathrm{S}}\sim z$ to leading-order, while $\partial_{z}u^{\mathrm{S}}$ decays exponentially for deep water waves.
@@ 2.1 Coherent structures in rotating and wave-averaged turbulence
To illustrate the similarity of rotating turbulence and turbulence beneath surface waves we conduct large eddy simulations of equation (1) with implicit grid-scale dissipation in a unit cube with three choices for $\bm{\Omega}$,
@@
Following previous isotropic homogeneous turbulence studies (e.g., Orszag & Patterson (1972)), we produce an initial condition for the three simulations by conducting a preliminary simulation initialized with kinetic energy spectrum
@@
The initial simulation is run for a duration $t_{0}$ until $t=0$, defined as the time when the mean-square vorticity has decayed to $\omega_{\text{rms}}(t=0)=10$. The velocity field is then saved to disk to be used as an initial condition in subsequent runs starting from $t=0$. The simulations are conducted with Oceananigans (Ramadhan et al., 2020; Wagner et al., 2025), which discretizes (1) with a finite volume method. Scripts that reproduce simulations in this paper are stored on GitHub; see the Data availability statement.
We focus first on the evolution of the relative vorticity $\bm{\omega}$ defined in equation (6). Figure 1 shows vorticity components for the three cases after $t=1000$ time units: figure 1(a) shows vertical vorticity $\zeta$, while while figures 1(b) and (c) show the horizontal vorticity $\eta$. Figure 1(a) and (b) for rotating and wave-affected turbulence, respectively, both exhibit the formation of coherent structures and relatively greater vorticity levels than the unorganized, small amplitude isotropic vorticity in figure 1(c). Figure 2 is similar, except that figure 2(a) shows $\zeta$ in rotating turbulence in the $xy$-plane, while figures 2(b) and 2(c) show $\eta$ in the $xz$ plane for wave-averaged and isotropic turbulence, respectively.
@@ 2.2 Zonation and the analogy of wave-averaged turbulence with beta-plane turbulence
Note that turbulence beneath shallow water waves is homeomorphic to $\beta$-plane turbulence Rhines (1975) — with $\bm{\Omega}_{\text{$\beta$-plane}}=\beta y\hskip 1.00006pt\bm{\hat{z}}$ — modulo on the $xz$ plane instead of on the $xy$ plane and with the Stokes drift curvature $\partial_{z}^{2}u^{\mathrm{S}}$ playing the role of $\beta$. One of the most striking results of this similarity is the propensity for turbulence beneath surface waves to develop “zonal jets” — coherent, alternating jets in the direction of surface wave propagation (and perpendicular to the pseudovorticity direction). This similarity allows us, therefore, to borrow intuition on structure formation in $\beta$-plane turbulence (for example, Huang & Robinson (1998); Farrell & Ioannou (2007); Srinivasan & Young (2012); Constantinou et al. (2014); for an overview see Constantinou (2015); Farrell & Ioannou (2019); Marston & Tobias (2023).)
@@
To illustrate this, we consider two additional cases with background vorticity
@@
$\bm{\Omega}_{\mathrm{deep}}=-\tfrac{1}{4}\mathrm{e}^{8(z-1)}\hskip 1.00006pt\bm{\hat{y}}\,.\qquad\text{and}\qquad\bm{\Omega}_{\mathrm{weak}}=-\tfrac{1}{8}z\hskip 1.00006pt\bm{\hat{y}}\,.$ (7)
@@
These and subsequent simulations use $384^{3}$ finite volume cells to reduce their computational expense for the purpose of performing longer simulations to $t=10^{4}$.
@@
Figure 3 shows time-series of the $y$-momentum $v$, which plays the role that vertical velocity plays in rotating turbulence. The right side of figure 3 shows vertical profiles of the horizontally-averaged $x$-momentum,
$U(z,t)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\int u\,\mathrm{d}x\,\mathrm{d}y\,.$ (8)
@@
These $U$-profiles exhibit the development of depth-alternating jets, which are most prominently exhibited by the zig-zagging structure in the medium waves case on the right side of figure 3. To our knowledge, depth-alternating jets have not been observed in shear-free wave-modified turbulence. Comparing the slices of $v$ for medium and weak waves reveals how the medium-strong waves induce a strong inverse cascade, more coherent vortices, and fewer small-scale motions than the weak waves.
@@
Figure 3: The evolution of cross-wave momentum $v$ in the $xz$-plane (6 panels on the left) and horizontally-averaged along-wave-momentum $u$ (2 panels on the right) at $t=40,400$ and for three wave fields: “deep” (left panel, red lines), “medium” (middle panel, orange lines), and “weak” (right panel, blue lines). Dashed lines in the top right plot show the Stokes drift profile (normalized) for each case. The “medium” waves case uses the background vorticity $\bm{\Omega}_{\text{waves}}$ in (4). Both the $u$ profiles and the light gray “zero lines” are spaced apart by the scale $\delta u=10^{-2}$. The $u$-profiles at right show the development of depth-alternating jets, including a counter-wave surface jet for all cases. The $v$-slices are similar between the three cases at early times, but at later times exhibit strong, localized wave-impacts in their respective regions of significant Stokes shear. Note that $v$ plays the role that vertical velocity plays in rotating turbulence. Note that the Rhines scale may be estimated as $\ell_{R}=2\pi\sqrt{U/\big{|}\partial_{z}^{2}u^{\mathrm{S}}\big{|}}$. With $\big{|}\partial_{z}^{2}u^{\mathrm{S}}\big{|}=\tfrac{1}{2}$ and $U\approx 10^{-2}$ for the medium waves case, for example, we estimate $\ell_{R}\approx 0.8$, consistent with the jet spacing apparent on the right panels.
@@ 3 Phenomenological model for the evolution of kinetic energy
Figure 4 plots time-series of the normalized kinetic energy $k(t)/k(t=0)$ for the three cases presented in figure 1, as well as three additional cases that use $\bm{\Omega}=Sz\hskip 1.00006pt\bm{\hat{y}}$ with $S=(1/2,1/8,1/16)$. Figure 4 illustrates another common feature to rotating turbulence and turbulence beneath surface waves: a suppression of the kinetic energy dissipation rate, such that at long times the kinetic energy levels off. We theorize that the suppression of dissipation is linked to the formation of coherent structures: rather than cascading to the grid scale, where dissipation can occur, kinetic energy accumulates in much larger-scale coherent structures. Similar to the dynamics of two-dimensional turbulence McWilliams (1984), the concentration of kinetic energy in coherent structures suppresses the forward cascade of kinetic energy.
Inspired by Bardina et al. (1985), we model the kinetic energy $k(t)$ with the phenomenological two-equation system
@@
The free parameter $a$ may be constrained by considering isotropic turbulence with $\Omega=0$. In this case we expect the turbulent kinetic energy to decay according to $k\sim t^{-6/5}$ (Saffman, 1967), which implies $\frac{\mathrm{d}}{\mathrm{d}t}k=-\frac{6}{5}\frac{k}{t}$, and thus in turn, via (10), leads to
$\epsilon=\frac{6}{5}\frac{k}{t}\,,\quad\text{so that}\quad\frac{\mathrm{d}}{\mathrm{d}t}\epsilon=-\frac{66}{25}\frac{k}{t^{2}}\,.$ (12)
@@
From (10)–(11) we deduce that the rate of change of the combination $\log\epsilon-a\log k$ is proportional to $b$, that is,
@@
$\frac{\epsilon}{\epsilon_{0}}=\left(\frac{k}{k_{0}}\right)^{a}\mathrm{e}^{-b\Omega t}\,,$ (14)
where $\epsilon_{0}$ and $k_{0}$ are the dissipation and kinetic energy at $t=0$ respectively. If we insert the expression for the dissipation $\epsilon(t)$ from (14) into (10) and integrate in time, we obtain
$\frac{k}{k_{0}}=\left[1+\frac{1}{nb\Omega}\frac{\epsilon_{0}}{k_{0}}\left(1-\mathrm{e}^{-b\Omega t}\right)\right]^{-n}\,,$ (15)
where $n=1/(a-1)=6/5$. (See also equation (10) in Bardina et al. (1985).) Taking the limit $\Omega\to 0$ (or $t\to 0$ with finite $\Omega$), we obtain the corresponding solution for isotropic turbulence,
@@
Figure 4: The decay of kinetic energy $k(t)$ in isotropic, rotating, and surface-wave-modulated turbulence. Solid lines show kinetic energy normalized by it’s initial value, $k/k_{0}$, computed from large eddy simulations. Dashed lines show $k/k_{0}$ given by (15), which solves the phenomenological two-equation system in (10)–(11). The initial dissipation rate $\epsilon_{0}$ in (15) is estimated from the numerical solution; specifically we evaluate (16) at $\Delta t$, substitute $k(t=\Delta t)$ from the numerical solution, and solve for $\epsilon_{0}$. We use $a=11/6$ and estimate $b$ from (18). For the rotating case, $\Omega=1/4$ yields $b=0.033$. For the four surface-wave-modulated cases we use $b=0.036$, where $\Omega\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\int\partial_{z}u^{\mathrm{S}}{\,\mathrm{d}}z=\left(1/4,1/8,1/16,1/32\right)$. Note that using $a=1.75$ along with commensurate adjustments to $b$ matches the simulation data even more closely.